OpenStudy (anonymous):

-2(5x-3(2x+(x+y))+3)

OpenStudy (anonymous):

make clear you question is it to solve !!!

OpenStudy (anonymous):

-2 (3 + 5 x - 3 (3 x + y))

OpenStudy (anonymous):

\[-2(5x-3(2x+(x+y))+3)=-2(5x-3(2x+x+y)+3)\] \[=-2(5x-3(3x+y)+3)\] \[=-2(4x-9x-3y+3)\] \[=-2(-5x-3y+3)\] \[=10x+6y+-6\] \[=2(5x+3y-3)\]